OpenStudy (anonymous):

Erik's age is 8 years less than one half his fathers, and Erik is also one third the age of his father. What is the age of Erik and his father?

OpenStudy (anonymous):

e = erik's age..... f = fathers age Erik's age is 8 years less than one half his fathers e = 1/2 f - 8 Erik is also one third the age of his father e = 1/3 f

OpenStudy (anonymous):

\[\{\text{erick}\to 16,f\to 48\} \]

OpenStudy (anonymous):

I need to show work done

OpenStudy (anonymous):

\[\text{erik}=\frac{f}{2}-8,\text{erik}=\frac{f}{3} \]\[\frac{f}{2}-8=\frac{f}{3} \]\[6 \left(-\frac{f}{3}+\frac{f}{2}-8\right)=0 \]\[f-48=0 \]\[f=48 \]\[\text{erik}=\frac{f}{3}=\frac{48}{3}=16 \]

OpenStudy (anonymous):

thank you so much robtobey

OpenStudy (anonymous):

You are welcome and thank you for the medal.